# Strict universality of the square-root law in price impact across stocks: a complete survey of the Tokyo stock exchange

### v2

@@ Abstract
Universal power laws have been scrutinised in physics and beyond, and a long-standing debate exists in econophysics regarding the strict universality of the nonlinear price impact, commonly referred to as the square-root law (SRL). The SRL posits that the average price impact II follows a power law with respect to transaction volume $Q$, such that $I(Q)\propto Q^{\delta}$ with $\delta\approx 1/2$. Some researchers argue that the exponent $\delta$ should be system-specific, without universality. Conversely, others contend that $\delta$ should be exactly $1/2$ for all stocks across all countries, implying universality. However, resolving this debate requires high-precision measurements of $\delta$ with errors of around $0.1$ across hundreds of stocks, which has been extremely challenging due to the scarcity of large microscopic datasets—those that enable tracking the trading behaviour of all individual accounts. Here we conclusively support the universality hypothesis of the SRL by a complete survey of all trading accounts for all liquid stocks on the Tokyo Stock Exchange (TSE) over eight years. Using this comprehensive microscopic dataset, we show that the exponent $\delta$ is equal to $1/2$ within statistical errors at both the individual stock level and the individual trader level. Additionally, we rejected two prominent models supporting the nonuniversality hypothesis: the Gabaix-Gopikrishnan-Plerou-Stanley and the Farmer-Gerig-Lillo-Waelbroeck models (Nature 2003, QJE 2006, and Quant. Finance 2013). Our work provides exceptionally high-precision evidence for the universality hypothesis in social science and could prove useful in evaluating the price impact by large investors—an important topic even among practitioners.
@@ Introduction.
The SRL is an empirical law that describes the nonlinear price response following large buying (selling) by institutional investors. Consider a trader buying (or selling) $Q$ units of stock by market orders—immediate decisions to buy or sell shares. Following this transaction, the market responds with a price movement $\Delta p$. For small $Q$, a linear response is expected, such that $I(Q):=\langle\epsilon\Delta p\mid Q\rangle\approx\lambda Q$, where $I(Q)$ is called the price impact [3], $\epsilon$ is the market order sign defined by $\epsilon=+1$ ($\epsilon=-1$) for buying (selling), and $\langle A\mid B\rangle$ represents the average of AA conditioned on $B$. For large $Q$, on the contrary, a nonlinear response is empirically observed (see Fig. 1),
@@
where $\langle A\rangle$ represents the cross-sectional average of AA on the TSE dataset across all stocks [^1] If we estimate $\delta$ by the nonlinear ordinary least squares on raw datapoints (without using binned averages), we obtain $\langle\delta_{\rm OLS}\rangle=0.474\pm 0.0012$. . This is the first main result of this article, suggesting that the exact SRL holds for more than two thousand data points. Also, Fig. 2(d) presents the histogram of the prefactor $c$. Furthermore, the crossover from the linear to square-root laws was observed consistently with Refs. [33, 4] (see SM [9]).
@@
For this purpose, we develop a simple statistical model with the exact SRL under order splitting and estimate the corresponding errorbar (see End Matter and the SM [9]): (i) We randomly shuffled the signs of metaorders, while keeping the physical times of market-order submissions identical to those in our TSE dataset. (ii) The price impacts were assumed to exactly obey the SRL with $\delta=1/2$, with random contributions. The model parameters were based on our dataset. (iii) We repeatedly generated price time series for 100 iterations and measured $\delta$ for all stocks (see Fig. 3(a) for a $\delta$’s histogram for one numerical iteration, corresponding to Fig. 2(c) for our TSE dataset). (iv) The average and standard deviation of $\delta$ were evaluated across all Monte Carlo trials as $\langle\!\langle\overline{\delta}\rangle\!\rangle$ and $\langle\!\langle\overline{\sigma_{\delta}}\rangle\!\rangle$. For our numerical simulation, $\langle\!\langle A\rangle\!\rangle$ represents the Monte Carlo average of AA over $N_{\rm MC}=100$ iterations, whereas $\overline{A}$ represents the average across stocks for one Monte Carlo iteration. This model generates plausible price time-series, consistent with diffusive price. Indeed, Ref. [36] develops a nonlinear price-impact model under order splitting, reminiscent of this statistical model, to show its consistency with various stylized facts.
@@ Acknowledgements.
We appreciate the careful reading of this manuscript by the staff of JPX Group. Also, we thank Jean-Phillipe Bouchaud for his motivating comments on our conference presentations and this manuscript. YS was supported by JSPS KAKENHI (Grant No. 24KJ1328). $K$ was supported by JSPS KAKENHI (Grant Nos. 21H01560, 22H01141, and 25K01450) and the JSPS Core-to-Core Program (Grant No. JPJSCCA20200001).
@@
The data supporting our results were provided by Japan Exchange (JPX) Group, Inc. JPX Group is a third-party commercial company and provided their dataset through a non-disclosure agreement with Kyoto University, strictly for academic purposes. This non-disclosure agreement imposes legal restrictions on data availability, and therefore, we cannot make the data publicly accessible without approval from JPX Group. However, upon reasonable request, $K$ will provide all the original data and programming code, subject to explicit approval by JPX Group.
@@
YS developed the program code and conducted the statistical analysis. $K$ designed and managed the project. Both YS and $K$ contributed to writing the manuscript and approved its final content.

### v3

@@ Abstract
Universal power laws have been scrutinised in physics and beyond, and a long-standing debate exists in econophysics regarding the strict universality of the nonlinear price impact, commonly referred to as the square-root law (SRL). The SRL posits that the average price impact $I$ follows a power law with respect to transaction volume $Q$, such that $I(Q)\propto Q^{\delta}$ with $\delta\approx 1/2$. Some researchers argue that the exponent $\delta$ should be system-specific, without universality. Conversely, others contend that $\delta$ should be exactly $1/2$ for all stocks across all countries, implying universality. However, resolving this debate requires high-precision measurements of $\delta$ with errors of around $0.1$ across hundreds of stocks, which has been extremely challenging due to the scarcity of large microscopic datasets—those that enable tracking the trading behaviour of all individual accounts. Here we conclusively support the universality hypothesis of the SRL by a complete survey of all trading accounts for all liquid stocks on the Tokyo Stock Exchange (TSE) over eight years. Using this comprehensive microscopic dataset, we show that the exponent $\delta$ is equal to $1/2$ within statistical errors at both the individual stock level and the individual trader level. Additionally, we rejected two prominent models supporting the nonuniversality hypothesis: the Gabaix-Gopikrishnan-Plerou-Stanley and the Farmer-Gerig-Lillo-Waelbroeck models (Nature 2003, QJE 2006, and Quant. Finance 2013). Our work provides exceptionally high-precision evidence for the universality hypothesis in social science and could prove useful in evaluating the price impact by large investors—an important topic even among practitioners.
@@ Introduction.
The SRL is an empirical law that describes the nonlinear price response following large buying (selling) by institutional investors. Consider a trader buying (or selling) $Q$ units of stock by market orders—immediate decisions to buy or sell shares. Following this transaction, the market responds with a price movement $\Delta p$. For small $Q$, a linear response is expected, such that $I(Q):=\langle\epsilon\Delta p\mid Q\rangle\approx\lambda Q$, where $I(Q)$ is called the price impact [3], $\epsilon$ is the market order sign defined by $\epsilon=+1$ ($\epsilon=-1$) for buying (selling), and $\langle A\mid B\rangle$ represents the average of $A$ conditioned on $B$. For large $Q$, on the contrary, a nonlinear response is empirically observed (see Fig. 1),
@@
where $\langle A\rangle$ represents the cross-sectional average of $A$ on the TSE dataset across all stocks [^1] If we estimate $\delta$ by the nonlinear ordinary least squares on raw datapoints (without using binned averages), we obtain $\langle\delta_{\rm OLS}\rangle=0.474\pm 0.0012$. . This is the first main result of this article, suggesting that the exact SRL holds for more than two thousand data points. Also, Fig. 2(d) presents the histogram of the prefactor $c$. Furthermore, the crossover from the linear to square-root laws was observed consistently with Refs. [33, 4] (see SM [9]).
@@
For this purpose, we develop a simple statistical model with the exact SRL under order splitting and estimate the corresponding errorbar (see End Matter and the SM [9]): (i) We randomly shuffled the signs of metaorders, while keeping the physical times of market-order submissions identical to those in our TSE dataset. (ii) The price impacts were assumed to exactly obey the SRL with $\delta=1/2$, with random contributions. The model parameters were based on our dataset. (iii) We repeatedly generated price time series for 100 iterations and measured $\delta$ for all stocks (see Fig. 3(a) for a $\delta$’s histogram for one numerical iteration, corresponding to Fig. 2(c) for our TSE dataset). (iv) The average and standard deviation of $\delta$ were evaluated across all Monte Carlo trials as $\langle\!\langle\overline{\delta}\rangle\!\rangle$ and $\langle\!\langle\overline{\sigma_{\delta}}\rangle\!\rangle$. For our numerical simulation, $\langle\!\langle A\rangle\!\rangle$ represents the Monte Carlo average of $A$ over $N_{\rm MC}=100$ iterations, whereas $\overline{A}$ represents the average across stocks for one Monte Carlo iteration. This model generates plausible price time-series, consistent with diffusive price. Indeed, Ref. [36] develops a nonlinear price-impact model under order splitting, reminiscent of this statistical model, to show its consistency with various stylized facts.
@@ Acknowledgements.
We appreciate the careful reading of this manuscript by the staff of JPX Group. Also, we thank Jean-Phillipe Bouchaud for his motivating comments on our conference presentations and this manuscript. YS was supported by JSPS KAKENHI (Grant No. 24KJ1328). KK was supported by JSPS KAKENHI (Grant Nos. 21H01560, 22H01141, and 25K01450) and the JSPS Core-to-Core Program (Grant No. JPJSCCA20200001).
@@
The data supporting our results were provided by Japan Exchange (JPX) Group, Inc. JPX Group is a third-party commercial company and provided their dataset through a non-disclosure agreement with Kyoto University, strictly for academic purposes. This non-disclosure agreement imposes legal restrictions on data availability, and therefore, we cannot make the data publicly accessible without approval from JPX Group. However, upon reasonable request, KK will provide all the original data and programming code, subject to explicit approval by JPX Group.
@@
YS developed the program code and conducted the statistical analysis. KK designed and managed the project. Both YS and KK contributed to writing the manuscript and approved its final content.